a 13-foot ladder is leaning against a house. if the base of the ladder is 5 feet from the house, how high up the side of the house will the ladder reach?
Answer by nerdybill(7384)

a 13-foot ladder is leaning against a house. if the base of the ladder is 5 feet from the house, how high up the side of the house will the ladder reach?
.
Applying Pythagorean theorem:
Let x = height ladder will reach
then
x^2 + 5^2 = 13^2
x^2 + 25 = 169
x^2 = 144
x = 12 feet